A newly MBA graduate received an $18000 signing bonus from her employer. She invested part of it in a credit union account earning 3% annual simple interest and the rest in utility stocks that suffered a 7% loss of value. The net income from both investments for the first year was only $90. How much of her bonus was originally placed in each investment?
Answer by checkley79(3341)

.03X-.07(18000-X)=90
.03X-1260+.07X=90
.10X=90+1260
.10X=1350
X=1350/.10
X=$13500 INVESTED @ 3%.
.03*13500-.07(18000-13500)=90
405-.07*4500=90
405-315=90
90=90